# Ortho-Radial Drawing in Near-Linear Time

### v2

@@ 1 Introduction
A plane graph is a planar graph $G=(V,E)$ with a combinatorial embedding $\mathcal{E}$. The combinatorial embedding $\mathcal{E}$ fixes a circular ordering $\mathcal{E}(v)$ of the edges incident to each vertex $v\in V$, specifying the counter-clockwise ordering of these edges surrounding $v$ in the drawing. An orthogonal drawing of a plane graph is a drawing of $G$ such that each edge is drawn as a sequence of horizontal and vertical line segments. For example, see Figure 1 for an orthogonal drawing of $K_{4}$ with four bends. Alternatively, an orthogonal drawing of $G$ can be seen as an embedding of $G$ into a grid such that the edges of $G$ correspond to internally disjoint paths in the grid. Orthogonal drawing is one of the most classical drawing styles studied in the field of graph drawing, and it has a wide range of applications, including VLSI circuit design [BL84, Val81], architectural floor plan design [LM81], and network visualization [BNT86, EGK+04, GJK+03, KDMW15].
@@
One of the most fundamental quality measures of orthogonal drawings is the number of bends. The bend minimization problem, which asks for an orthogonal drawing with the smallest number of bends, has been extensively studied over the past 40 years [CK12, DLV98, DLOP20, Sto80, Tam87, GT97]. In a seminal work, Tamassia [Tam87] introduced the topology-shape-metric framework to tackle the bend minimization problem. Tamassia showed that an orthogonal drawing can be described combinatorially by an orthogonal representation, which consists of an assignment of an angle in $\{90^{\circ},180^{\circ},270^{\circ},360^{\circ}\}$ to each corner and a designation of the outer face. In this paper, a corner is defined as a pair of edges incident to the same vertex that are consecutive in the given combinatorial embedding. Specifically, Tamassia [Tam87] showed that an orthogonal representation can be realized as an orthogonal drawing with zero bends if and only if the following two conditions are satisfied:
@@
An orthogonal representation is valid if it satisfies the above conditions (O1) and (O2). Given a valid orthogonal representation, an orthogonal drawing realizing the orthogonal representation can be computed in linear time [Hsu80, Tam87]. This result (shape $\rightarrow$ metric) allows us to reduce the task of finding a bend-minimized orthogonal drawing (topology $\rightarrow$ metric) to the conceptually much simpler task of finding a bend-minimized valid orthogonal representation (topology $\rightarrow$ shape).
By focusing on orthogonal representations, we may neglect certain geometric aspects of graph drawing such as edge lengths and vertex coordinates, making the task of algorithm design easier. In particular, given a fixed combinatorial embedding, the task of finding a bend-minimized orthogonal representation can be easily reduced to the computation of a minimum cost flow [Tam87]. Such a reduction to a flow computation is not easy to see if one thinks about orthogonal drawings directly without thinking about orthogonal representations.
@@ 1.1 Ortho-radial drawing
The study of ortho-radial drawing is motivated by its applications in network visualization [BBS21, FLW14, XCC22], particularly in the context of transit map layout [WNT+20]. Ortho-radial drawings are especially well-suited for visualizing metro systems with radial and circular routes. Examples of such drawings can be found in [Ove03, Rob05, Rob12].
@@
Given a plane graph, an ortho-radial representation is defined as an assignment of an angle to each corner together with a designation of the central face and the outer face. Barth, Niedermann, Rutter, and Wolf [BNRW23] showed that an ortho-radial representation can be realized as an ortho-radial drawing with zero bends if the following three conditions are satisfied:
@@
The definition of a strictly monotone cycle is technical and depends on the choice of the reference edge $e^{\star}$, so we defer its formal definition to a subsequent section. The reference edge $e^{\star}$ is an edge in the contour of the outer face and is required to lie on the outermost circular arc used in an ortho-radial drawing. Informally, a strictly monotone cycle has a structure that is like a loop of ascending stairs or a loop of descending stairs, so a strictly monotone cycle cannot be drawn. The necessity of (R1)–(R3) is intuitive to see. The more challenging and interesting part of the proof in [BNRW23] is to show that these three conditions are actually sufficient.
@@ 1.2 Previous methods
The journal paper [BNRW23] is a combination of two works [BNRW17, NRW19]. In the first work [BNRW17], the proof that conditions (R1)–(R3) are necessary and sufficient is only existential in that it does not yield efficient algorithms to check the validity of a given ortho-radial representation and to construct an ortho-radial drawing without bends realizing a given ortho-radial representation.
Checking (R1) and (R2) can be done in linear time in a straightforward manner. The difficult part is to design an efficient algorithm to check (R3). The most naive approach of examining all cycles costs exponential time. The second work [NRW19] addressed this gap by showing an $O(n^{2})$-time algorithm to decide whether a strictly monotone cycle exists for a given reference edge $e^{\star}$, where $n$ is the number of vertices in the input graph. They also show an $O(n^{2})$-time algorithm to construct an ortho-radial drawing without bends, for any given ortho-radial representation with a reference edge $e^{\star}$ that does not contain a strictly monotone cycle.
@@
The main idea behind the proof in the first work [BNRW17] is a reduction to the easier case where each regular face is rectangular. For this case, they provided a proof that conditions (R1)–(R3) are necessary and sufficient, and they also provided an efficient drawing algorithm via a reduction to a flow computation given that (R1)–(R3) are satisfied.
For any given ortho-radial representation with $n$ vertices, it is possible to add $O(n)$ additional edges to turn it into an ortho-radial representation where each regular face is rectangular. A major difficulty in the proof of [BNRW17] is that they need to ensure that the addition of the edges preserves not only (R1) and (R2) but also (R3). The lack of an efficient algorithm to check whether (R3) is satisfied is precisely the reason that the proof of [BNRW17] does not immediately lead to a polynomial-time algorithm.
@@
The above issue was addressed in the second work [NRW19]. They provided an $O(n^{2})$-time algorithm to find a strictly monotone cycle if one exists, given a fixed choice of the reference edge $e^{\star}$. This immediately leads to an $O(n^{2})$-time algorithm to decide whether a given ortho-radial representation, with a fixed reference edge $e^{\star}$, admits an ortho-radial drawing. Moreover, combining this $O(n^{2})$-time algorithm with the proof of [BNRW17] discussed above yields an $O(n^{4})$-time drawing algorithm. The time complexity is due to the fact that $O(n)$ edge additions are needed for rectangulation, for each edge addition there are $O(n)$ candidate reference edges to consider, and to test the feasibility of each candidate edge they need to run the $O(n^{2})$-time algorithm to test whether the edge addition creates a strictly monotone cycle.
@@
Using further structural insights on the augmentation process of [BNRW17], the time complexity of the above $O(n^{4})$-time drawing algorithm can be lowered to $O(n^{2})$ [NRW19]. The reason for the quadratic time complexity is that for each of the $O(n)$ edge additions, a left-first DFS starting from the newly added edge is needed to test whether the addition of this edge creates a strictly monotone cycle.
@@ 1.3 Our new method
For both validity testing (checking whether a given angle assignment induces a strictly monotone cycle) and drawing (finding a geometric embedding realizing a given ortho-radial representation), the two algorithms in [BNRW23] naturally cost $O(n^{2})$ time, as they both require performing left-first DFS $O(n)$ times.
@@
The core of our method is the notion of a good sequence, which we briefly explain below. An ortho-radial representation satisfying (R1) and (R2), with a fixed reference edge $e^{\star}$, determines whether an edge $e$ is a vertical edge (i.e., $e$ is drawn as a segment of a straight line passing through the origin) or horizontal (i.e., $e$ is drawn as a circular arc of some circle centered on the origin). Let $E_{\mathsf{h}}$ denote the set of horizontal edges, oriented in the clockwise direction, and let $\mathcal{S}_{\mathsf{h}}$ denote the set of connected components induced by $E_{\mathsf{h}}$. Note that each component $S\in\mathcal{S}_{\mathsf{h}}$ is either a path or a cycle.
@@
A major difference between our method and the approach based on rectangulation in [BNRW23] is that the cost for adding a new virtual edge is only $O(\log n)$ in our algorithm. As we will later see, in our algorithm, in order to identify new virtual edges to be added, we only need to do some simple local checks such as calculating the sum of angles, and there is no need to do a full left-first DFS to test whether a newly added edge creates a strictly monotone cycle.
@@
While we show a nearly linear-time algorithm for the (shape $\rightarrow$ metric)-step (i.e., from ortho-radial representations to ortho-radial drawings), essentially nothing is known about the (topology $\rightarrow$ shape)-step (from planar graphs to ortho-radial representations). While the task of finding a bend-minimized orthogonal representation of a given plane graph can be easily reduced to the computation of a minimum cost flow [Tam87], such a reduction does not apply to ortho-radial representations, as network flows do not work well with the notion of strictly monotone cycles. It remains an open question as to whether a bend-minimized ortho-radial representation of a plane graph can be computed in polynomial time.
@@ 1.4 Related work
Orthogonal drawing is a central topic in graph drawing, see [DG13] for a survey. The bend minimization problem for orthogonal drawings of planar graphs of maximum degree $4$ without a fixed combinatorial embedding is NP-hard [FHH+93, GT97]. If the combinatorial embedding is fixed, the topology-shape-metric framework of Tamassia [Tam87] reduces the bend minimization problem to a min-cost flow computation. The algorithm of Tamassia [Tam87] costs $O(n^{2}\log n)$ time. The time complexity was later improved to $O\left(n^{7/4}\sqrt{\log n}\right)$ [GT97] and then to $O\left(n^{3/2}\log n\right)$ [CK12]. A recent $O(n\operatorname{poly}\log n)$-time planar min-cost flow algorithm [DGG+22] implies that the bend minimization problem can be solved in $O(n\operatorname{poly}\log n)$ time if the combinatorial embedding is fixed.
If the combinatorial embedding is not fixed, the NP-hardness result of [FHH+93, GT97] can be bypassed if the first bend on each edge does not incur any cost [BRW16] or if we restrict ourselves to some special class of planar graphs. In particular, for planar graphs with maximum degree 3, it was shown that the bend-minimization can be solved in polynomial time [DLV98]. After a series of improvements [CY17, DLOP20, DLP18], we now know that a bend-minimized orthogonal drawing of a planar graph with maximum degree $3$ can be computed in $O(n)$ time [DLOP20].
The topology-shape-metric framework [Tam87] is not only useful in bend minimization, but it is also, implicitly or explicitly, behind the graph drawing algorithms for essentially all computational problems in orthogonal drawing and its variants, such as morphing orthogonal drawings [BLPS13], allowing vertices with degree greater than 4 [DBDPP99, KM98, PT00], restricting the direction of edges [DLP14, DFMM14], drawing cluster graphs [BCFW04], and drawing dynamic graphs [BW98].
The study of ortho-radial drawing by Barth, Niedermann, Rutter, and Wolf [BNRW23] extended the topology-shape-metric framework [Tam87] to accommodate cylindrical grids. Before the work [BNRW23], a combinatorial characterization of drawable ortho-radial representation was only known for paths, cycles, and theta graphs [HHT09], and for the special case where the graph is 3-regular and each regular face in the ortho-radial representation is a rectangle [HHMT10].
@@ 1.5 Organization
In Section 2, we discuss the basic graph terminology used in this paper, review some results in the previous work [BNRW23], and state our main theorems. In Section 3, we introduce the notion of a good sequence and show that its existence implies a simple ortho-radial drawing algorithm. In Section 4, we present a greedy algorithm that adds virtual edges to a given ortho-radial representation with a fixed reference edge so that a good sequence that covers the entire graph exists and can be computed efficiently. In Section 5, we show that a strictly monotone cycle, which certifies the non-existence of a drawing, exists and can be computed efficiently if the greedy algorithm fails. In Section 6, we show that our results can be extended to the setting where the reference edge is not fixed at the cost of an extra logarithmic factor in the time complexity. In Section 7, we justify our assumption that the input graph is simple and biconnected by showing a reduction from any graph to a biconnected simple graph. We conclude in Section 8 with discussions on possible future directions.
@@ 2 Preliminaries
Throughout the paper, let $G=(V,E)$ be a planar graph of maximum degree at most $4$ with a fixed combinatorial embedding $\mathcal{E}$ in the sense that, for each vertex $v\in V$, a circular ordering $\mathcal{E}(v)$ of its incident edges is given to specify the counter-clockwise ordering of these edges surrounding $v$ in a planar embedding. As we will discuss in Section 7, we may assume that the input graph $G$ is simple and biconnected. In this section, we introduce some basic graph terminology and review some results from Barth, Niedermann, Rutter, and Wolf [BNRW23].
@@
Unless otherwise stated, all edges, paths, and cycles are assumed to be directed. We write $\overline{e}$, $\overline{P}$, and $\overline{C}$ to denote the reversal of an edge $e$, a path $P$, and a cycle $C$, respectively. We allow paths and cycles to have repeated vertices and edges. We say that a path or a cycle is simple if it does not have repeated vertices. Following [BNRW23], we say that a path or a cycle is crossing-free if it satisfies the following conditions:
@@
It was shown in [BNRW23] that $(\mathcal{R},e^{\star})$ is drawable if and only if the ortho-radial representation $\mathcal{R}$ satisfies (R1) and (R2) with the reference edge $e^{\star}$ does not contain a strictly monotone cycle. Since it is straightforward to test whether (R1) and (R2) are satisfied in linear time, from now on, unless otherwise stated, we assume that (R1) and (R2) are satisfied for the ortho-radial representation $\mathcal{R}$ under consideration.
@@
For any single-edge path $P$, we define $\mathsf{rotation}(P)=0$. For any crossing-free path $P$ of length more than $2$, we define $\mathsf{rotation}(P)$ to be the sum of the combinatorial rotations of all length-2 subpaths of $P$. Similarly, for any cycle $C$ of length more than $2$, we define $\mathsf{rotation}(C)$ to be the sum of the combinatorial rotations of all length-2 subpaths of $C$. Same as [BNRW23], based on this notion, we may restate condition (R2) as follows.
@@
The equivalence between the new and the old definitions of (R2) stems from the fact that a $90^{\circ}$ left turn corresponds to an angle of $270^{\circ}$. If $F$ is a regular face with $k$ corners, then in the original definition of (R2), it is required that the sum $s$ of angles around $F$ is $s=(k-2)\cdot 180^{\circ}$. Since the facial cycle $C_{F}$ traverses the contour of $F$ in the clockwise direction, the number of $90^{\circ}$ right turns minus the number of $90^{\circ}$ left turns must be exactly $4$. Therefore, $s=(k-2)\cdot 180^{\circ}$ is the same as $\mathsf{rotation}(C_{F})=4$, as each $90^{\circ}$ right turn contributes $+1$ and each $90^{\circ}$ left turn contributes $-1$ in the calculation of $\mathsf{rotation}(C_{F})$.
@@
Let $C$ be a simple cycle oriented in such a way that the outer face $F_{\mathsf{o}}$ lies in its exterior. Following [BNRW23], we say that $C$ is essential if the central face $F_{\mathsf{c}}$ is in the interior of $C$. Otherwise we say that $C$ is non-essential. The following lemma was proved in [BNRW23].
Lemma 2.1 ([BNRW23]).
Suppose (R1) and (R2) are satisfied. Let $C$ be a simple cycle oriented in such a way that the outer face $F_{\mathsf{o}}$ lies in its exterior, then the combinatorial rotation of $C$ satisfies the following condition.
@@
Following [BNRW23], for any two edges $e=(u,v)$ and $e^{\prime}=(x,y)$, we say that a crossing-free path $P$ is a reference path for $e$ and $e^{\prime}$ if $P$ starts at $u$ or $v$ and ends at $x$ or $y$ such that $P$ does not contain any of the edges in $\{e,\overline{e},e^{\prime},\overline{e^{\prime}}\}$. Given a reference path $P$ for $e=(u,v)$ and $e^{\prime}=(x,y)$, we define the combinatorial direction of $e^{\prime}$ with respect to $e$ and $P$ as follows.
@@
Consider the reference edge $e=(v_{14},v_{1})$ in the ortho-radial representation of Figure 3. We measure the direction of $e^{\prime}=(v_{8},v_{9})$ from $e$ with different choices of the reference path $P$. If $P=(v_{1},v_{2},v_{9})$, then $\mathsf{direction}(e,P,e^{\prime})=\mathsf{rotation}(e\circ P\circ\overline{e^{\prime}})-2=-1$. If $P=(v_{14},v_{10},v_{9})$, then we also have $\mathsf{direction}(e,P,e^{\prime})=\mathsf{rotation}(\overline{e}\circ P\circ\overline{e^{\prime}})=-1$. If we select $P=(v_{1},v_{2},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8})$, then we get a different value of $\mathsf{direction}(e,P,e^{\prime})=\mathsf{rotation}({e}\circ P\circ{e^{\prime}})=3$. As we will discuss later, $\mathsf{direction}(e,P,e^{\prime})\operatorname{mod}4$ is invariant under the choice of $P$ [BNRW23].
@@
We say that $e$ is a vertical edge if $e$ points towards north or south. Otherwise, we say that $e$ is a horizontal edge. We argue that as long as (R1) and (R2) are satisfied, the direction of any edge $e$ is uniquely determined by the ortho-radial representation $\mathcal{R}$ and the reference edge $e^{\star}$.

### v3

@@ 1 Introduction
A plane graph is a planar graph $G=(V,E)$ with a combinatorial embedding $\mathcal{E}$. The combinatorial embedding $\mathcal{E}$ fixes a circular ordering $\mathcal{E}(v)$ of the edges incident to each vertex $v\in V$, specifying the counter-clockwise ordering of these edges surrounding $v$ in the drawing. An orthogonal drawing of a plane graph is a drawing of $G$ such that each edge is drawn as a sequence of horizontal and vertical line segments. For example, see Figure 1 for an orthogonal drawing of $K_{4}$ with four bends. Alternatively, an orthogonal drawing of $G$ can be seen as an embedding of $G$ into a grid such that the edges of $G$ correspond to internally disjoint paths in the grid. Orthogonal drawing is one of the most classical drawing styles studied in the field of graph drawing, and it has a wide range of applications, including VLSI circuit design [bhatt1984framework, valiant1981universality], architectural floor plan design [LIGGETT1981277], and network visualization [batini1986layout, eiglsperger2004automatic, gutwenger2003new, kieffer2015hola].
@@
One of the most fundamental quality measures of orthogonal drawings is the number of bends. The bend minimization problem, which asks for an orthogonal drawing with the smallest number of bends, has been extensively studied over the past 40 years [CK12, BLV98, didimo2020optimal, GT96, storer1980node, tamassia1987embedding]. In a seminal work, Tamassia [tamassia1987embedding] introduced the topology-shape-metric framework to tackle the bend minimization problem. Tamassia showed that an orthogonal drawing can be described combinatorially by an orthogonal representation, which consists of an assignment of an angle in $\{90^{\circ},180^{\circ},270^{\circ},360^{\circ}\}$ to each corner and a designation of the outer face. In this paper, a corner is defined as a pair of edges incident to the same vertex that are consecutive in the given combinatorial embedding. Specifically, Tamassia [tamassia1987embedding] showed that an orthogonal representation can be realized as an orthogonal drawing with zero bends if and only if the following two conditions are satisfied:
@@
An orthogonal representation is valid if it satisfies the above conditions 1 and 2. Given a valid orthogonal representation, an orthogonal drawing realizing the orthogonal representation can be computed in linear time [hsueh1980symbolic, tamassia1987embedding]. This result (shape $\rightarrow$ metric) allows us to reduce the task of finding a bend-minimized orthogonal drawing (topology $\rightarrow$ metric) to the conceptually much simpler task of finding a bend-minimized valid orthogonal representation (topology $\rightarrow$ shape).
By focusing on orthogonal representations, we may neglect certain geometric aspects of graph drawing such as edge lengths and vertex coordinates, making the task of algorithm design easier. In particular, given a fixed combinatorial embedding, the task of finding a bend-minimized orthogonal representation can be easily reduced to the computation of a minimum cost flow [tamassia1987embedding]. Such a reduction to a flow computation is not easy to see if one thinks about orthogonal drawings directly without thinking about orthogonal representations.
@@ 1.1 Ortho-radial drawing
The study of ortho-radial drawing is motivated by its applications in network visualization [bast2021metro, fink2014concentric, xu2022automated], particularly in the context of transit map layout [wu2020survey]. Ortho-radial drawings are especially well-suited for visualizing metro systems with radial and circular routes. Examples of such drawings can be found in [ovenden2003metro, roberts2005underground, roberts2012underground].
@@
Given a plane graph, an ortho-radial representation is defined as an assignment of an angle to each corner together with a designation of the central face and the outer face. Barth, Niedermann, Rutter, and Wolf [barth2023topology] showed that an ortho-radial representation can be realized as an ortho-radial drawing with zero bends if the following three conditions are satisfied:
@@
The definition of a strictly monotone cycle is technical and depends on the choice of the reference edge $e^{\star}$, so we defer its formal definition to a subsequent section. The reference edge $e^{\star}$ is an edge in the contour of the outer face and is required to lie on the outermost circular arc used in an ortho-radial drawing. Informally, a strictly monotone cycle has a structure that is like a loop of ascending stairs or a loop of descending stairs, so a strictly monotone cycle cannot be drawn. The necessity of 13 is intuitive to see. The more challenging and interesting part of the proof in [barth2023topology] is to show that these three conditions are actually sufficient.
@@ 1.2 Previous methods
The journal paper [barth2023topology] is a combination of two works [barth2017, niedermann19]. In the first work [barth2017], the proof that conditions 13 are necessary and sufficient is only existential in that it does not yield efficient algorithms to check the validity of a given ortho-radial representation and to construct an ortho-radial drawing without bends realizing a given ortho-radial representation.
Checking 1 and 2 can be done in linear time in a straightforward manner. The difficult part is to design an efficient algorithm to check 3. The most naive approach of examining all cycles costs exponential time. The second work [niedermann19] addressed this gap by showing an $O(n^{2})$-time algorithm to decide whether a strictly monotone cycle exists for a given reference edge $e^{\star}$, where $n$ is the number of vertices in the input graph. They also show an $O(n^{2})$-time algorithm to construct an ortho-radial drawing without bends, for any given ortho-radial representation with a reference edge $e^{\star}$ that does not contain a strictly monotone cycle.
@@
The main idea behind the proof in the first work [barth2017] is a reduction to the easier case where each regular face is rectangular. For this case, they provided a proof that conditions 13 are necessary and sufficient, and they also provided an efficient drawing algorithm via a reduction to a flow computation given that 13 are satisfied.
For any given ortho-radial representation with $n$ vertices, it is possible to add $O(n)$ additional edges to turn it into an ortho-radial representation where each regular face is rectangular. A major difficulty in the proof of [barth2017] is that they need to ensure that the addition of the edges preserves not only 1 and 2 but also 3. The lack of an efficient algorithm to check whether 3 is satisfied is precisely the reason that the proof of [barth2017] does not immediately lead to a polynomial-time algorithm.
@@
The above issue was addressed in the second work [niedermann19]. They provided an $O(n^{2})$-time algorithm to find a strictly monotone cycle if one exists, given a fixed choice of the reference edge $e^{\star}$. This immediately leads to an $O(n^{2})$-time algorithm to decide whether a given ortho-radial representation, with a fixed reference edge $e^{\star}$, admits an ortho-radial drawing. Moreover, combining this $O(n^{2})$-time algorithm with the proof of [barth2017] discussed above yields an $O(n^{4})$-time drawing algorithm. The time complexity is due to the fact that $O(n)$ edge additions are needed for rectangulation, for each edge addition there are $O(n)$ candidate reference edges to consider, and to test the feasibility of each candidate edge they need to run the $O(n^{2})$-time algorithm to test whether the edge addition creates a strictly monotone cycle.
@@
Using further structural insights on the augmentation process of [barth2017], the time complexity of the above $O(n^{4})$-time drawing algorithm can be lowered to $O(n^{2})$ [niedermann19]. The reason for the quadratic time complexity is that for each of the $O(n)$ edge additions, a left-first DFS starting from the newly added edge is needed to test whether the addition of this edge creates a strictly monotone cycle.
@@ 1.3 Our new method
For both validity testing (checking whether a given angle assignment induces a strictly monotone cycle) and drawing (finding a geometric embedding realizing a given ortho-radial representation), the two algorithms in [barth2023topology] naturally cost $O(n^{2})$ time, as they both require performing left-first DFS $O(n)$ times.
@@
The core of our method is the notion of a good sequence, which we briefly explain below. An ortho-radial representation satisfying 1 and 2, with a fixed reference edge $e^{\star}$, determines whether an edge $e$ is a vertical edge (i.e., $e$ is drawn as a segment of a straight line passing through the origin) or horizontal (i.e., $e$ is drawn as a circular arc of some circle centered on the origin). Let $E_{\mathsf{h}}$ denote the set of horizontal edges, oriented in the clockwise direction, and let $\mathcal{S}_{\mathsf{h}}$ denote the set of connected components induced by $E_{\mathsf{h}}$. Note that each component $S\in\mathcal{S}_{\mathsf{h}}$ is either a path or a cycle.
@@
A major difference between our method and the approach based on rectangulation in [barth2023topology] is that the cost for adding a new virtual edge is only $O(\log n)$ in our algorithm. As we will later see, in our algorithm, in order to identify new virtual edges to be added, we only need to do some simple local checks such as calculating the sum of angles, and there is no need to do a full left-first DFS to test whether a newly added edge creates a strictly monotone cycle.
@@
While we show a nearly linear-time algorithm for the (shape $\rightarrow$ metric)-step (i.e., from ortho-radial representations to ortho-radial drawings), essentially nothing is known about the (topology $\rightarrow$ shape)-step (from planar graphs to ortho-radial representations). While the task of finding a bend-minimized orthogonal representation of a given plane graph can be easily reduced to the computation of a minimum cost flow [tamassia1987embedding], such a reduction does not apply to ortho-radial representations, as network flows do not work well with the notion of strictly monotone cycles. It remains an open question whether a bend-minimized ortho-radial representation of a plane graph can be computed in polynomial time.
@@ 1.4 Related work
Orthogonal drawing is a central topic in graph drawing, see [DG13] for a survey. The bend minimization problem for orthogonal drawings of planar graphs of maximum degree $4$ without a fixed combinatorial embedding is NP-hard [formann1993drawing, GT96]. If the combinatorial embedding is fixed, the topology-shape-metric framework of Tamassia [tamassia1987embedding] reduces the bend minimization problem to a min-cost flow computation. The algorithm of Tamassia [tamassia1987embedding] costs $O(n^{2}\log n)$ time. The time complexity was later improved to $O\left(n^{7/4}\sqrt{\log n}\right)$ [GT96] and then to $O\left(n^{3/2}\log n\right)$ [CK12]. A recent $O(n\operatorname{poly}\log n)$-time planar min-cost flow algorithm [dong2022nested] implies that the bend minimization problem can be solved in $O(n\operatorname{poly}\log n)$ time if the combinatorial embedding is fixed.
If the combinatorial embedding is not fixed, the NP-hardness result of [formann1993drawing, GT96] can be bypassed if the first bend on each edge does not incur any cost [BRW12] or if we restrict ourselves to some special class of planar graphs. In particular, for planar graphs with maximum degree 3, it was shown that the bend-minimization can be solved in polynomial time [BLV98]. After a series of improvements [chang2017bend, didimo2020optimal, didimo2018bend], we now know that a bend-minimized orthogonal drawing of a planar graph with maximum degree $3$ can be computed in $O(n)$ time [didimo2020optimal].
The topology-shape-metric framework [tamassia1987embedding] is not only useful in bend minimization, but it is also, implicitly or explicitly, behind the graph drawing algorithms for essentially all computational problems in orthogonal drawing and its variants, such as morphing orthogonal drawings [biedl2013morphing], allowing vertices with degree greater than 4 [BDPP99, KM98, papakostas2000efficient], restricting the direction of edges [didimo2014complexity, durocher2014drawing], drawing cluster graphs [brandes2004draw], and drawing dynamic graphs [brandes1998dynamic].
The study of ortho-radial drawing by Barth, Niedermann, Rutter, and Wolf [barth2023topology] extended the topology-shape-metric framework [tamassia1987embedding] to accommodate cylindrical grids. Before the work [barth2023topology], a combinatorial characterization of drawable ortho-radial representation was only known for paths, cycles, and theta graphs [hasheminezhad2009ortho], and for the special case where the graph is 3-regular and each regular face in the ortho-radial representation is a rectangle [hasheminezhad2010rectangular].
@@ 1.5 Organization
In Section 2, we discuss the basic graph terminology used in this paper, review some results in the previous work [barth2023topology], and state our main theorems. In Section 3, we introduce the notion of a good sequence and show that its existence implies a simple ortho-radial drawing algorithm. In Section 4, we present a greedy algorithm that adds virtual edges to a given ortho-radial representation with a fixed reference edge so that a good sequence that covers the entire graph exists and can be computed efficiently. In Section 5, we show that a strictly monotone cycle, which certifies the non-existence of a drawing, exists and can be computed efficiently if the greedy algorithm fails. In Section 6, we show that our results can be extended to the setting where the reference edge is not fixed at the cost of an extra logarithmic factor in the time complexity. In Section 7, we justify our assumption that the input graph is simple and biconnected by showing a reduction from any graph to a biconnected simple graph. We conclude in Section 8 with discussions on possible future directions.
@@ 2 Preliminaries
Throughout the paper, let $G=(V,E)$ be a planar graph of maximum degree at most $4$ with a fixed combinatorial embedding $\mathcal{E}$ in the sense that, for each vertex $v\in V$, a circular ordering $\mathcal{E}(v)$ of its incident edges is given to specify the counter-clockwise ordering of these edges surrounding $v$ in a planar embedding. As we will discuss in Section 7, we may assume that the input graph $G$ is simple and biconnected. In this section, we introduce some basic graph terminology and review some results from Barth, Niedermann, Rutter, and Wolf [barth2023topology].
@@
Unless otherwise stated, all edges, paths, and cycles are assumed to be directed. We write $\overline{e}$, $\overline{P}$, and $\overline{C}$ to denote the reversal of an edge $e$, a path $P$, and a cycle $C$, respectively. We allow paths and cycles to have repeated vertices and edges. We say that a path or a cycle is simple if it does not have repeated vertices. Following [barth2023topology], we say that a path or a cycle is crossing-free if it satisfies the following conditions:
@@
It was shown in [barth2023topology] that $(\mathcal{R},e^{\star})$ is drawable if and only if the ortho-radial representation $\mathcal{R}$ satisfies 1 and 2 with the reference edge $e^{\star}$ does not contain a strictly monotone cycle. Since it is straightforward to test whether 1 and 2 are satisfied in linear time, from now on, unless otherwise stated, we assume that 1 and 2 are satisfied for the ortho-radial representation $\mathcal{R}$ under consideration.
@@
For any single-edge path $P$, we define $\mathsf{rotation}(P)=0$. For any crossing-free path $P$ of length more than $2$, we define $\mathsf{rotation}(P)$ to be the sum of the combinatorial rotations of all length-2 subpaths of $P$. Similarly, for any cycle $C$ of length more than $2$, we define $\mathsf{rotation}(C)$ to be the sum of the combinatorial rotations of all length-2 subpaths of $C$. Same as [barth2023topology], based on this notion, we may restate condition 2 as follows.
@@
The equivalence between the new and the old definitions of 2 stems from the fact that a $90^{\circ}$ left turn corresponds to an angle of $270^{\circ}$. If $F$ is a regular face with $k$ corners, then in the original definition of 2, it is required that the sum $s$ of angles around $F$ is $s=(k-2)\cdot 180^{\circ}$. Since the facial cycle $C_{F}$ traverses the contour of $F$ in the clockwise direction, the number of $90^{\circ}$ right turns minus the number of $90^{\circ}$ left turns must be exactly $4$. Therefore, $s=(k-2)\cdot 180^{\circ}$ is the same as $\mathsf{rotation}(C_{F})=4$, as each $90^{\circ}$ right turn contributes $+1$ and each $90^{\circ}$ left turn contributes $-1$ in the calculation of $\mathsf{rotation}(C_{F})$.
@@
Let $C$ be a simple cycle oriented in such a way that the outer face $F_{\mathsf{o}}$ lies in its exterior. Following [barth2023topology], we say that $C$ is essential if the central face $F_{\mathsf{c}}$ is in the interior of $C$. Otherwise we say that $C$ is non-essential. The following lemma was proved in [barth2023topology].
Lemma 2.1 ([barth2023topology]).
Suppose 1 and 2 are satisfied. Let $C$ be a simple cycle oriented in such a way that the outer face $F_{\mathsf{o}}$ lies in its exterior, then the combinatorial rotation of $C$ satisfies the following condition.
@@
Following [barth2023topology], for any two edges $e=(u,v)$ and $e^{\prime}=(x,y)$, we say that a crossing-free path $P$ is a reference path for $e$ and $e^{\prime}$ if $P$ starts at $u$ or $v$ and ends at $x$ or $y$ such that $P$ does not contain any of the edges in $\{e,\overline{e},e^{\prime},\overline{e^{\prime}}\}$. Given a reference path $P$ for $e=(u,v)$ and $e^{\prime}=(x,y)$, we define the combinatorial direction of $e^{\prime}$ with respect to $e$ and $P$ as follows.
@@
Consider the reference edge $e=(v_{14},v_{1})$ in the ortho-radial representation of Figure 3. We measure the direction of $e^{\prime}=(v_{8},v_{9})$ from $e$ with different choices of the reference path $P$. If $P=(v_{1},v_{2},v_{9})$, then $\mathsf{direction}(e,P,e^{\prime})=\mathsf{rotation}(e\circ P\circ\overline{e^{\prime}})-2=-1$. If $P=(v_{14},v_{10},v_{9})$, then we also have $\mathsf{direction}(e,P,e^{\prime})=\mathsf{rotation}(\overline{e}\circ P\circ\overline{e^{\prime}})=-1$. If we select $P=(v_{1},v_{2},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8})$, then we get a different value of $\mathsf{direction}(e,P,e^{\prime})=\mathsf{rotation}({e}\circ P\circ{e^{\prime}})=3$. As we will discuss later, $\mathsf{direction}(e,P,e^{\prime})\operatorname{mod}4$ is invariant under the choice of $P$ [barth2023topology].
@@
We say that $e$ is a vertical edge if $e$ points towards north or south. Otherwise, we say that $e$ is a horizontal edge. We argue that as long as 1 and 2 are satisfied, the direction of any edge $e$ is uniquely determined by the ortho-radial representation $\mathcal{R}$ and the reference edge $e^{\star}$.